OpenStudy (anonymous):

how do i answer 2x-3=sqrroot30-7x

OpenStudy (anonymous):

2x-3=\[\sqrt{30-7x}\]

OpenStudy (mrnood):

is this \[2x-3=\sqrt{30-7x}\] OR \[2x-3=\sqrt{30} - 7x\]

OpenStudy (anonymous):

the first one

OpenStudy (mrnood):

ok soz - I saw your answer after I posted! square both sides of the equation

OpenStudy (anonymous):

its okay

OpenStudy (mrnood):

\[(2x-3)^{2}=(\sqrt{30-7x})^{2}\]

OpenStudy (anonymous):

so the square root cancels

OpenStudy (mrnood):

yes

OpenStudy (anonymous):

and 2x-3 turns into 4x^2-12x+9

OpenStudy (mrnood):

yes Then collect all terms onto 1 side (other side =0) You then have a quadratic

OpenStudy (anonymous):

okay so ima have 0=4x^2+5x-21

OpenStudy (mrnood):

not sure that is correct 4x^2-12x+7x+9-30

OpenStudy (anonymous):

so 4x^2-5x-21 ???

OpenStudy (mrnood):

yes It doesn't factorise - so use formula

OpenStudy (anonymous):

what do i do next??

OpenStudy (mrnood):

use quadratic formula to solve the equation 4x^2-5x-21=0 (ax^2+bx+c=0) x=(-b +/- sqrt (b^2-4ac))/2a

OpenStudy (anonymous):

i forgot wAT to do after this part 5x+\[\sqrt{361}\] over 8